2. How many molecules are in 3.87 moles of CCl 3. How many oxygen atoms are in 2.50 moles of CH20


mole calculations, with an explanation on the set up
Answer #1

2)
1 mole of CCl4 has 6.02*1023 molecule
3.87 mole of CCl4 has (6.02*1023*3.87) molecule
number of molecule of CCl4 = 6.02*1023*3.87
number of molecule of CCl4 = 23.3*1023

3)
1 mole of C6H12O6 has 6.02*1023 molecule
2.5 mole of C6H12O6 has (6.02*1023*2.5) molecule
number of molecule of C6H12O6 = 6.02*1023*2.5
number of molecule of C6H12O6 = 15.1*1023

1 molecule of C6H12O6 has 6 O atom
15.1*1023 molecule of C6H12O6 has (6*15.1*1023) O atom
number of O atom = 6*15.1*1023
number of O atom = 90.6*1023